OpenStudy (anonymous):

How would u get rid of the pi in this equation: 25pi(r)(r+2pi(r)^2

OpenStudy (anonymous):

it should be equal to something ?

OpenStudy (anonymous):

all this equals the surface area of a given cylinder

OpenStudy (anonymous):

can u type r without brackets and only use where neceessary

OpenStudy (anonymous):

okay, the original was . S=2pir(r+25/pir^2)

OpenStudy (anonymous):

pi is getting cancells directly as one pi is in Numerator and other is in denominator!!

OpenStudy (anonymous):

ooo k. i see that. but for the answer they have 2pir^2+50r^-1. they brought up the pir^2 from the denominator to make 2pir(r+25pir^-2)

OpenStudy (anonymous):

u need to type the question correctly earlier u typed it differently please upload the questionor attach file

OpenStudy (anonymous):

Originially It yped it this way, i see now that as a typo I wrote 2pir^-2 instead of 25. The Original : S= 2pir(r+25/pir^2) They simplified it so that now it's S= 2pir^2+50r^-1 I understand the 2pir was distributed nto in the parenthesis, but I do not understand why the pi was cancelled in the simplified version.

OpenStudy (anonymous):

\[ S= 2r \pi(r+ \frac{ 25 }{ \pi^2 })\] Is this the correct version? If so, then \[ S= (2r \pi)r + (2r \pi)(\frac{ 25 }{ \pi^2 })\] notice that the pi will cancel at the back, leaving, \[S=2\pi r ^2 + \frac{ 50r }{ \pi}\]

OpenStudy (anonymous):

its just simple multiplication results in cancellation of pi!!

OpenStudy (anonymous):

ok, thank you

OpenStudy (anonymous):

You're welcome!